Review the material in Chapter 20 and respond to the following:

Discuss the modern quantity theory and the liquidity preference theory.
Answer #1

Ans ) Liquidity preference theory suggest that investors will demand higher rate of interest for the medium term and long term securities because they prefer cash as the most suitable liquid asset and will invest in long term or medium term securities only if they have more premium of interest as the liquidity is compromised in this case of long term investment as there is a larger risk associated with that.And the liquidity is preferred for speculation and precautionery measure .Modern quantity theory predicts that the demand for money not only depends upom the risk and return but is alsp depended on the various various types of assets an individual can hold.The choice based on the wealth and hence the demand for money is function of wealth.That can be understood by the fisher formula which is as follows :The formula for money demand as per fisher is as follows : M *V = P*T

where M = stock of money in coin , notes and bank deposit and V = rate of circulation or the frequency of transaction.P = measure of price index like consumer price index and T = The total volume of money transaction in the economy during a year.Hence stock of money multiplied by rate of circulation should equals the volume of money transaction multiplied by price index.The other important measure is friedman formula , which states that , Y = C + I +G + ( X-M)

here C = consumption expenditure

I = Investment expenditure.

G = government spending

X - M = exports - imports .

Here Y is income in the economy in dollar terms and ia the supply function and on the right is the demand function = C +I + G +( X-M)